OpenStudy (jackellyn):

Use the definition to find the laplace transform of the following function from the graph

OpenStudy (jackellyn):

 slope=y/t=15/5=3 y intercept =0 y=mt+b y=3t

OpenStudy (anonymous):

So you're trying to convert y(t) to Y(s)?

OpenStudy (jackellyn):

Trying to find the Laplace transform, our notation might be different.

OpenStudy (anonymous):

The laplace transform of 3t is 3/s^2 if that's what you're trying to find..

OpenStudy (jackellyn):

Could you show me your work using the definition? I have been trying to figure out all the little details for hours.

OpenStudy (anonymous):

Ok, lets just take the laplace transform of T, and we can just multiply 3 to it at the end. Here's the definition for the laplace transform of T. http://gyazo.com/272d2975504bfd1c2fdc47e174c955f8.png

OpenStudy (jackellyn):

Sorry, I can't see anything when I click